A television network is deciding whether or not to give its newest television show a spot during prime happen, it will have to move one of its most viewed shows to another slot. The show they would rather watch. The network will keep its current lineup of shows unless the majority of the customers want to the new show. The network recelves 827 responses, of which 428 indicate that they would like to see the new show in the lineup. (You may find it useful to reference the appropriate table: z table or t table viewing time at night. For this to network conducts a survey asking its viewers which a. Set up the hypotheses to test if the television network should give its newest television show a spot during prime viewing time at night. b-1. Calculate the value of the test statistic. (Round intermediate calculations to at least 4 decimal places and decimal places.) final answer to 2 Test statistic
b-2. Find the p-value p-value < 0.01 0.01 s p-value<0.025 0.025 s p-value < 0.05 0.05 s p-value <0.10 Op-value 2 0.10 c. At a 0.01, what should the television network do? Reject Ho, the television network should not keep its current lineup. Do not reject Ho, the television network should keep its current lineup.
Answer #1

a). H0 : p < 0.5

    Ha : p > 0.5 (Claim) Viewers needs to vote in majority.

b).

b-1.) Using sample data, we calculate the standard deviation (σ) and compute the z-score test statistic (z).

σ = sqrt[ P * ( 1 - P ) / n ] = sqrt [(0.5 * 0.5) / 827]

σ = 0.0174

z = (p - P) / σ = (0.52 - .50)/0.0174 = 1.1494

Test statistic = 1.1494

The P-Value is 0.125196.
The result is not significant at p < 0.01.

b-2). p-value > 0.10

c). Since the P-value (0.125196) is greater than the significance level (0.01), we cannot reject the null hypothesis.

Do not reject Ho, the television network should keep its current lineup.